OpenStudy (amazingbread12):

Which equation could be used to solve the problem? Troy typed 375 word in 15 minutes. How many words per minute did he type (w)? A. w = 375 words ÷ 15 minutes B. w = 15 minutes ÷ 375 words C. w = 375 words + 15 minutes D. w = 375 words • 15 minutes

OpenStudy (brooklynsshaw):

It cannot be B otherwise the answer would be negative, and Troy would not be able to type in a negative amount of time

OpenStudy (brooklynsshaw):

Excuse me, i meant a decimal

OpenStudy (anonymous):

And it can't be C. because you don't add when you do this, you divide. And the answer could be a decimal.

OpenStudy (xmissalycatx):

You would want to divide so that you can determine how many words per EACH minute that he typed..

OpenStudy (brooklynsshaw):

^

OpenStudy (anonymous):

So, the answer is A.? Right?

OpenStudy (brooklynsshaw):

I agree

OpenStudy (xmissalycatx):

Correct!

OpenStudy (brooklynsshaw):

Good work :)

OpenStudy (xmissalycatx):

I'll medal you @brooklynsshaw

OpenStudy (anonymous):

Thank you so much :) Would you mind if you could help with a few more plz?

OpenStudy (brooklynsshaw):

Thank you darling !! :) @xMissAlyCatx

OpenStudy (brooklynsshaw):

No problem, what are your other questions?

OpenStudy (anonymous):

Which equation could be used to solve the problem? Oscar averaged 4 yards per carry in the football game. If he carried the ball 16 times, how many yards did he gain (y)? A. y = 4 + 16 B. y = 4 ÷ 16 C. y = 16 ÷ 4 D. y = 4 • 16

OpenStudy (brooklynsshaw):

Multiply 4 by 16

OpenStudy (xmissalycatx):

This will be a multiplication situation because you want to know how many all together he got. SO what @brooklynsshaw said just now..

OpenStudy (brooklynsshaw):

y = 4 x 16

OpenStudy (brooklynsshaw):

Exactly @xMissAlyCatx :)

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

There are 192 children at camp and they are evenly divided into 8 groups. How many children are in each group?

OpenStudy (xmissalycatx):

This will be division because you want to know how many kids PER group.

OpenStudy (anonymous):

ok, so 8 divided by 192? Or 192 divided by 8?

OpenStudy (xmissalycatx):

192 divided by 8. If you do a small divided by a large number it will be a fraction/decimal.

OpenStudy (brooklynsshaw):

The answer will be the product of 192 divided by 8

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

Ms. Huffman bought 15 cases of juice boxes for the town picnic. There were 24 juice boxes in each case. How many juice boxes did she buy?

OpenStudy (xmissalycatx):

You must multiply these two numbers to get how many there are all together.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the unit rate for 540 calories in 90 minutes? https://static.k12.com/bank_packages/files/media/mathml_17b4dce2b0056171cab9ff10d031da08f2f1f4ea_1.gif

OpenStudy (xmissalycatx):